OpenStudy (anonymous):

for which value of c will roots of 2x^2-4x+=0 be real rational and equal?

OpenStudy (anonymous):

Your question seems to be missing something, which I'm going to attempt to fill in. 2x^2-4x+c=0 (2x-4)(x-2)=0 This would make both the roots x=2. Therefore, c is(-4)(-2)=8.

OpenStudy (anonymous):

nice job ranger!